OpenStudy (anonymous):

f-1(3) when f(x) = 2 x minus 7, all over 3

OpenStudy (anonymous):

\(\sf \large f(x) = \frac{2x-7}{3}\) to get the inverse: 1. Let f(x)= y, so it will be like : \(\sf \large y = \frac{2x-7}{3}\) 2. Switch the places of x and y variables: \(\sf \large x= \frac{2y-7}{3}\) 3. Solve the equation in terms of y. Tell me what will you get. 4.Change y into \(\sf f(x)^{-1}\) 5. Solve \(\sf f(3)^{-1}\) by substituting x=3. Tell me also what will you get in this part.

OpenStudy (anonymous):

How do I solve it? I'm confused

OpenStudy (anonymous):

hmm.. let's start with the first step. you just have to replace f(x) with y to make things easier :) is this part clear?

OpenStudy (anonymous):

Yes I understand that part its solving \[2y - 7 \over 3\] that I don't get how to do.

OpenStudy (anonymous):

ahh okay great. so you have: 

OpenStudy (anonymous):

well i tired it and got \[3x + 7 \over 2 \]

OpenStudy (anonymous):

which would = y

OpenStudy (anonymous):

alright let's see :)  That's perfect! good job :D so you're done with step 3. I'll do step 4 for you. Replace y into f(x)^-1 \(\sf \Large f(x)^{-1}= \frac {3x + 7}{ 2}\) Now, the final step :) subsitute x= 3 in the equation and simplify!